OpenStudy (anonymous):

I don't get how to do this problem. The tables below show solutions for two linear equations. If the two equations make up a system of linear equations, in which quadrant is the solution? X Y X Y -10 9 -4 -10 -7 3 -3 -3 -4 -4 -2 4 A. Quadrant I B. Quadrant II C. Quadrant III D. Quadrant IV I know that the first table is in quadrant II and the second one has two dots in the III Quadrant but one is in the IV Quadrant.

OpenStudy (amistre64):

you have to define a slope for each table

OpenStudy (anonymous):

How do you do that?

OpenStudy (amistre64):

a slope is the difference between 2 points lets subtract one point from another in one table: -10 9 -(-7 3) ------ -3 6 ; and stac k y/x to define a numerical value for slope. 6/-3 = -2 the first table has a slope of -2

OpenStudy (amistre64):

the second table is easier since it moves the xs by 1; the slope is just read then as the difference in the ys: -3 -4 = -7 for its slope

OpenStudy (anonymous):

So you minus -10 and -9 with -7 and 3?

OpenStudy (amistre64):

you subtract the components of one slope from the other, yes y1 - y2 ------ = slope is the general formula for it x1 - x2

OpenStudy (amistre64):

subtract the components of one POINT from the other ... have to use the right words :)

OpenStudy (anonymous):

Okay so you would subtract -10 and -9 with each other. Do you divide at all?

OpenStudy (amistre64):

yes; slope is defined as the ratio of the change (the difference) in y as compared to the x

OpenStudy (anonymous):

So divide after you would subtract?

OpenStudy (amistre64):

step 1: subtract one point from the other step 2: stack y over x. y/x

OpenStudy (anonymous):

Okay, so that gives you the slope. What do you do with the slope to solve the answer?

OpenStudy (amistre64):

(-10, 9) -( -7, 3) (-10, 9) ( 7, -3) ---------- -3 , 6 slope = y/x: 6/-3 = -2

OpenStudy (amistre64):

you use the slopes, and given point, to construct line equations to compare with

OpenStudy (anonymous):

Oh so you make a line and see what quadrant their in. Then that gives you the answer?

OpenStudy (amistre64):

i have a concern tho: -7 3 --4 --4 -------- -3 7 this last point on the first table is NOT in line with the others

OpenStudy (amistre64):

the place the 2 lines cross gives you the answer; its a process that combines all the skills that youve been working on together to get you to determine a solution to a more complicated problem.

OpenStudy (anonymous):

How is it not in line. It doesn't have to be straight.

OpenStudy (amistre64):

lines, by default, are linear .. they are straight from end to end

OpenStudy (amistre64):

the points you have provided for table 1 are not in a line with each other

OpenStudy (amistre64):

check to see that you havent typoed an error into it

OpenStudy (anonymous):

Oh it's suppose to be -4 and -3...my bad.

OpenStudy (amistre64):

to double check X Y X Y -10 9 -4 -10 +3 -6 +1 +7 ------- ------- -7 3 -3 -3 +3 -6 +1 +7 ------ ------- -4 -3 -2 4 thats good then :) we have a slope of -6/3 and 7/1

OpenStudy (amistre64):

now im going to build equations for the lines using the slopes AND a known point from the line. y = -2x+2(-7)+3 y = 7x-7(-2)+4 = -2x -14 +3 = 7x +14+4 = -2x -11 =7x +18 now, where these 2 lines cross gives me a point in some quadrant: we can solve this "system of equations" to determine where they meet

OpenStudy (anonymous):

Okay so what are the lines? -2, -11 and 7, 18?

OpenStudy (amistre64):

the lines are what are formed by the equations; -2,-11 is not a line; it would be a point someplace on the graph .... the lines are: y = -2x-11 and y = 7x+18

OpenStudy (anonymous):

So they are points. Isn't that only for one line though?

OpenStudy (amistre64):

youll have to tell me where you got lost at. whats the last thing that made sense?

OpenStudy (anonymous):

You said that y = -2 x -11 and y = 7x + 18 but isn't that only one line. You need two lines to find out where they meet. One of the dots in in quadrant II and the other is in Quadrant I.

OpenStudy (anonymous):

Not meet... where they cross each other.

OpenStudy (amistre64):

y= one line ; and y= the other line those are the 2 equations; one for each line the dots given are just there to aid you in forming the lines; the lines themselves are made up of infinitely many points.

OpenStudy (amistre64):

 at the moment our lines look something like this; we just have to narrow down where they actually cross each other at

OpenStudy (anonymous):

Ok so how do we figure that out? Do we solve the equation?

OpenStudy (amistre64):

we solve the "system of equations". either by elimination, or substitution.

OpenStudy (amistre64):

our 2 equations are: y = -2x-11 y = 7x+18 which I havent test to make sure but I am fairly confident they are good :)

OpenStudy (phi):

btw, the question only asks for which quadrant the solution lies in. I would answer this question by a quick sketch.

OpenStudy (amistre64):

i got nothing reliable to sketch with; so to be sure im going the "system of equations" route

OpenStudy (anonymous):

So what would we substitute for x?

OpenStudy (amistre64):

id rather substitute for "y" :) since we have 2 choices ... y = -2x-11 so: y = 7x+18 becomes: -2x-11 = 7x+18 and solve for x

OpenStudy (amistre64):

-29 = 9x -29/9 = x use this to confirm a "y" value y = -2(-29/9) - 11 = 58/9 - 11 = (58 - 99)/9 i see is already going to be negative so id say ... (-,-) is quadrant 3

OpenStudy (amistre64):

and the wolf says its good :) http://www.wolframalpha.com/input/?i=y%3D-2x-11+and+y%3D7x%2B18

OpenStudy (anonymous):

So the answer is Quadrant III?

OpenStudy (amistre64):

that is what i deduce from it , yes

OpenStudy (amistre64):

2 1 3 4 yeah, q3

OpenStudy (anonymous):

Okay thank you, you really helped me :-)

OpenStudy (amistre64):

youre welcome :)